760 J of heat is used to increase the temperature of 400g of copper block from 30 ℃ to 35 ℃. What is the specific heat capacity of the copper block?


1
Expert's answer
2021-06-22T11:06:03-0400

"Q=cm(t_2-t_1)\to c=\frac{Q}{m(t_2-t_1)}=\frac{760}{0.4\cdot(35-30)}=380\ (J/(kg\cdot K))" . Answer